

********************************************************************
UARS WEEKLY STATUS REPORT - 27 NOV 1998
********************************************************************
***********
SOLSTICE
***********
NOTHING RECEIVED.
***********
SUSIM
***********
NOTHING RECEIVD.
***********
PEM
***********
NOTHING RECEIVED.
***********
HALOE
***********
NOTHING RECEIVED.
***********
MLS
***********
NOTHING RECEIVED.
***********
HRDI
***********
NOTHING RECEIVED.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING November 27, 1998
WINDII is preparing to include PMC observations during its
southward looking yaw period. This requires a 0.5 degree
roll-down because of the higher orbit altitude in the
southern hemisphere. We appreciate the support for that.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF notified all PI representatives regarding unrecoverable data gaps
for UARS DAYS 2624 and 2629 due to problems encountered at POCC.
The CDHF completed several Y2K time shift tests on a subset of the UARS
VMS cluster on November 24th. The system returned to current day without
incident. Mutinet problems were expected and were fixed.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HAL (4.80) PROD (L1) (1875R-1875C=0L)
30 days HAL (4.80) PROD (L2-L3) (62R-62C=0L)
80 days HAL (4.80) PROD (L3) (1874R-1429C=445L)
47 days SOL (4.62) PROD (L2-L3) (2313R-2111C=202L)
103 days SUS (7.00) PROD (L1) (2370R-507C=1863L)
3 days CLA (7.00) ENG (L1-L3) (24R-15C=9L)
1 day SEN (5.00) PROD (1R-1C=0L)
The CDHF installed the Y2K ECO patch (Number VAXY2K01_U2055) on
CDHF2, CDHF5, and UARSMS which addresses Y2K VMS operating system
issues.
DAAC data created for September 1998 has been loaded into the DAAC Transfer
Tables for HALOE, HRDI, NMC, PEM, and UKMO.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 13 November 1998 (Orbit 39209, GMT Day 318) through 20
November 1998 (Orbit 39298, GMT Day 324).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 318/2346 39223 Off
WINDII 318/2346 39223 Off
PEM 319/0318 39225 Off
PEM 320/0507 39241 On
SUSIM 321/0834 39258 Off
SUSIM 322/0007 39268 On
PEM 323/2337 39298 Axis Off
The beta angle ranged between 45.5 degrees to 73.1 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 77.0
degrees on 22 November 1998 (DOY 326).
Flight Operations continues to work with the PEM instrument to fine tune
the operational command sequences associated with the OBC Power Monitor
day-only operational capability. Realtime PEM day-only troubleshooting was
suspended when PEM went to full-time operations this week. Testing will
continue when PEM returns to day-only mode.
A power deficiency was experienced this week (21 November) due to
instrument scheduling confusion. Instrument operations were adjusted to
assure adequate battery state-of-charge.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.5 degrees C. Additional battery performance data
for each battery on 13 November 1998 and 20 November 1998 are provided in
Appendix D.
The clock error ranged from -6.2 to 11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 06
% HALOE 01
% PEM 04
% SUSIM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 06
Three (3) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496917
Transponder B: 98/300 - 2287.498980
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/323 show the observatory data loss to be
21 hours, 14 minutes, 49.268 seconds (an increase of 2 minutes and 56.704
seconds since last report period). This is a 0.0338 percent data loss
which equals a 99.9662 percent data capture for the mission.
The increase of 2 minutes and 56.704 seconds occurred over seven (7) days:
7 1.0 minute 7.840 seconds on Day 317 due to hits in data (IP);
7 1.664 seconds on Day 318 due to hits in data (IP);
7 3.936 seconds on Day 319 due to hits in data (IP);
7 1.216 seconds on Day 320 due to hits in data and insufficient overlap;
7 1.0 minute 37.056 seconds on Day 321 due to insufficient overlap;
7 1.056 seconds on Day 322 due to hits in data; and
7 3.936 seconds on Day 323 due to hits in data (IP).
Telemetry and Command Computer (TAC) RcrashesS have diminished this report
period. The cause of the RcrashesS has been identified by MSOCC and
corrective action implemented. The problem (as reported by MSOCC - Tom
Marsh) seems to have been produced with the data bursts experienced with
the IP network. The old MSS MDM systems provided data at a very periodic
rate; whereas, the IP network provides data in bursts. Many hardware tests
were performed to determine why the computer would stop executing
instructions. No hardware problem could be found. A conclusion was drawn
that it may be a timing problem. NASCOM was requested to add a two byte
delay between NASCOM blocks. This change seemed to improve performance.
NASCOM was then requested on 09 November to change the clock rate of the
data line for Q-channel data. The following week, no additional TAC
problems were reported.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The UARS FOT prepared and tested the Spacecraft plan for the Leonid
encounter on the UTTS. During the encounter, the Fixed Star Trackers
remained ON and operational based on a re-evaluation that turn off was a
greater risk. The Earth Sensor 2 was turned on and the spacecraft was
selected for earth pointing safehold (if the Onboard computer failed and
safehold was executed). UARS was configured on 17 November in Orbit 39262
at 321/1452 UTC and returned to normal configuration in Orbit 39268 at
322/0007 UTC. No anomalies have been observed on UARS.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - FOT attempted to participate in a Y2K UPS/NCC-
98 test on 19 November. However, due to some configuration problems with
the NCC-98 system, the test has been postponed. The test was supposed to
allow UPS users to perform scheduling operations with NCC (via the NCC-98
and the back-up UPS system) while a simulated clock roll-over from 1999 to
2000 occurred..
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - FOT received one (of two expected) Y2K 8-
week S/C Orbital Events file from FDF this week. The Orbital Events file
contained the previously missing SC_YAWMAN event. With the receipt of this
file, the last remaining Y2K test function (#12) for the 2000-to-2001 test
cases (#10 and 11) has been completed. The FOT also received a similar
file for the 1999-to-2000 roll-over, but the file contained orbit numbers
inconsistent with the previously generated 8-week PSAT file. FDF will
correct this orbit number problem and transmit a new file next week.
Flight Dynamics Facility (FDF) - see CMS paragraph.
IP NETWORK TESTING - Complete
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
FOT continues to prepare the Latch valve command sequences and to plan the
testing with the UTTS simulator in preparation for the simulation with
hardware. A test ATC load containing two PDB activation commands was built
this week. The purpose of the test ATC load was to simulate S/C thruster
firing sequences via PDBs with the UTTS. The test proved to be successful.
Engineering Division (Dennis Asato) advises that their engineer will not be
available until late January 1999 ( a change from the end of November 1998)
due to higher priority tasks. However, the UARS FOT is now ready to
support the tests.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing assistance on both sustaining engineering
and re-engineering activities. FOT was contacted by re-engineering (Steve
Liu) regarding the daily transfer of the FOTUs CMS loads to the UARS re-
engineering team. Apparently, re-engineering is having some minor problems
with interpreting the load formats. In addition, re-engineering is
experiencing a difference between the CMS generated checksums for the loads
and the checksums generated by the new ground system after the load
transfer is completed.
Testing of the Trending System v.1.1 release (CM-0009) continued. More
DR's and ER's are being generated.
MEETINGS
The next battery telecon is re-scheduled for 30 November 1998 in the power
branch conference room.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 13
December 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 20 November 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39223 318/2346 98-446 HALOE Power OFF
39223 318/2346 98-447 WINDII Power OFF
39225 319/0318 98-448 PEM Disable PMON Grp #10
39226 319/0501 98-449 OBC Clk Adj fr 0.00 to 20.6 msec/day
39241 320/0507 98-450 PEM Power ON
39258 321/0834 98-451 SUSIM Power OFF
39262 321/1452 98-452 All Pre-Leonid Configuration
39268 322/0007 98-453 SUSIM Power ON
39269 322/0148 98-454 All Post-Leonid Re-configuration
39279 322/1815 98-455 OBC Clk Adj fr 20.6 to 41.2 msec/day
39298 323/2337 98-456 PEM AXIS Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 20 November 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HATBAT Y-HI Bolometer Area Temp 39222
SOSGTEVLTG Y-HI G Tube Voltage 39238
HRPTSBARTEMP Y-LO Telescope Barrel Temp 39279-83
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 20 November 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39226 319/18:31 (15 Nov 98) from 00.0 to 20.6 msec/day
39279 322/05:55 (18 Nov 98) from 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 November 1998
13 Nov 1998 - GMT Day 317 Beta = 45.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +100.8/+84.0 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +33.6/-16.8 +5.15/+3.87 +28.8/-15.2 25.8 1.026 20.8
3 +145.6/+11.2 +4.19/+2.92 +28.0/-15.2 25.8 1.023 20.4
20 Nov 1998 - GMT Day 324 Beta = 73.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +100.8/+78.4 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +16.8/-22.4 +3.87/+1.97 +18.8/-15.6 27.0 1.034 16.0
3 +78.4/+11.2 +2.60/+1.34 +18.4/-15.2 26.9 1.033 15.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 20 November 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT 39228 319/0824 14171 Generic Late Acquisition
98-103 FOT 39239 320/0131 20377 Negative acquisition due
to an IP problem at STGT.
98-104 FOT 39249 320/1811 20383 IP lines misconfigured
98-105 FOT 39277 322/1340 20389 Operator Error
98-105 FOT 39294 323/1820 N/A OBC Dump errors
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-103 FOT 39239 320/01:31 20377 NEG ACQ-IP PROB
98-104 FOT 39249 320/18:11 20383 IP LINE CONF
98-105 FOT 39277 322/13:40 20389 OPN ERROR
98-106 FOT 39292 323/1500 TBL 33&34 DMP ERR
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT Nov 98 See List 14170 GENERIC LATE ACQ
98-107 FOT 39308 324/16:40 TAC 2 PROB
98-108 FOT 39309 324/17:23 TAC PROBLEMS
SEVNTFW47